OpenStudy (anonymous):

\[2x ^{2} + 11x -21 \] Factor of the polynomial

OpenStudy (anonymous):

(x+7)(2x-3)

OpenStudy (anonymous):

i factored, or u could use the quadratic formula.....

OpenStudy (anonymous):

Could you show me the formula's? I wanna be able to work and understand the problem.

OpenStudy (anonymous):

\[x = -b +/- radical b2 - 4ac/2a\]

OpenStudy (anonymous):

when you do so much factoring in school, it becomes like multiplication.

OpenStudy (anonymous):

whatever

OpenStudy (anonymous):

I am confused! xD How do i solve it where do i start?

OpenStudy (anonymous):

do u know the quadratic formula?

OpenStudy (anonymous):

no, sorry I don't this is all new to me

OpenStudy (anonymous):

ok, let me show you

OpenStudy (anonymous):

thanks! I appreciate it a lot!

OpenStudy (anonymous):

x = -b+/-radicalb2 - 4ac/2a, a,b,c,x2,x, integer.....

OpenStudy (anonymous):

the problem is 2x2 + 11x - 21, let a = 2, b = 11, c = -21.....

OpenStudy (anonymous):

the plug in the formula, let x = -11 =/- radical121 - 4 times 2 times - 21/2 times 2.....

OpenStudy (anonymous):

simplify, eventually it becomes, x = -7, 3/2.....

OpenStudy (anonymous):

righting this in factor form means making individual equations!!!!!

OpenStudy (anonymous):

easy.....

OpenStudy (anonymous):

1. x = -7, (x+7) = 0, 2. x = 3/2, (2x - 3) = 0.....

OpenStudy (anonymous):

He doesnt want to solve it......:/

OpenStudy (anonymous):

thus, (x+7)(2x-3) = 2x2 + 11x - 21.....

OpenStudy (anonymous):

there is no reason to show him quadratic formula. You have to learn both ways of solving for x so you can't avoid factoring by just showing the quadratic formula.

OpenStudy (anonymous):

can u leave ryanl, your violating the rules of open study, so contribute, or leave

OpenStudy (anonymous):

the quadratic formula is the easiest way for starters since all your doing is plugging in numbers.....

OpenStudy (anonymous):